5. Ysabelle deposits P8000 in one simple interest account and P2000 in a second simple interest account. The interest rate on the P8000 account is 2% more than the rate on the P2000 account. If the total yearly amount of interest on the two accounts is P578, find the interest rate on each account.
6. If in Problem 5, the total yearly amount of interest earned on the P8000 account is P459 more than the interest earned on the P2000 account, find the interest rate on each account.

Answer by mananth(16949)

8000 - (x+2)%
2000 - x%
...
0.01x*2000 + 0.01(x+2)*8000= 578
20x+80x+160=578
100x=578-160
100x=418
x= 4.18% - 8000
6.18% - 2000
...
The next part is similar type.
....
m.ananth@hotmail.com